write an equation in the slope-intercept form of the line passing through the points (3,-12)(-1,4)
Answer by tutorcecilia(2144)

First, find the slope of the line by using the formula for a slope:
Slope = m = (y1-y2)/(x1-x2) [plug-in the values and solve]
Slope = m=(-12-4)/[3-(-1)]=-16/4=-4
.
So, the slope = -4.
.
Now, use the formula for the slope-intercept equation of a line:
y=mx+b [plug-in the values and solve for the b-term]
.
-12=(-4)(3)+b [pick any point. I used (3, -12)]
-12=-12+b
-12+12=b
0=b [plug the b-term back into the formula along with the m-term]
.
y=mx+b
y=-4x+0
or y=-4x [the equation of the line in slope-intercept form]
.
Check by plugging one of the points back into the equation:
(-1, 4)
y=mx+b
4=(-4)(-1)+0
4=4 [checks out]